# Sound and Complete Neuro-symbolic Reasoning with LLM-Grounded Interpretations

### v1

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in natural language understanding and generation, but they exhibit problems with logical consistency in the output they generate. How can we harness LLMs’ broad-coverage parametric knowledge in formal reasoning despite their inconsistency? We present a method for directly integrating an LLM into the interpretation function of the formal semantics for a paraconsistent logic. We provide experimental evidence for the feasibility of the method by evaluating the function using datasets created from several short-form factuality benchmarks. Unlike prior work, our method offers a theoretical framework for neuro-symbolic reasoning that leverages an LLM’s knowledge while preserving the underlying logic’s soundness and completeness properties.
@@ 1 Introduction
LLMs encode vast parametric knowledge (Petroni et al., 2019) yet remain inconsistent and incomplete (Cheng et al., 2025), hindering their use as knowledge bases. Current approaches to logical reasoning with LLMs rely primarily on prompting strategies and external symbolic solvers (Wei et al., 2022; Cheng et al., 2025) but lack formal frameworks for managing the inherent inconsistency and incompleteness of LLM knowledge. Meanwhile, paraconsistent logics (Priest et al., 2025) are multi-valued non-classical logics designed to handle inconsistent information without logical explosion, where any contradiction would make everything provable. Belnap computers (Belnap, 1977a, b) are theoretical constructions described by Nuel Belnap to model contexts in which machines are responsible for reasoning in the face of incomplete or inconsistent information. This raises a natural question: can paraconsistent logic and Belnap computers provide a principled framework for using LLMs as knowledge bases despite their inconsistency and incompleteness?
We propose a Belnap computer whose external knowledge source is an LLM judge. LLM judges scale the process of evaluating LLM output for factuality given a short- or long-form question answering task, returning truth valuations for one or more statements in the LLM’s output (Li et al., 2024). In our approach, an LLM judge responds to a query for the valuation of an atomic formula in the context of a paraconsistent reasoner with a generalized truth value (Shramko and Wansing, 2025, 2011). Generalized truth values allow the LLM judge to provide information to the reasoner not only about the degree of truth of an atomic formula given an LLM’s parametric knowledge, but also with respect to the degree of knowledge the LLM has about the formula, in terms of whether it can verify or refute it.
@@
Section 2 discusses related work, and in the following sections we make three key contributions. Section 3 defines (i) a bilateral factuality evaluation function that provides useful information beyond that provided by current LLM judge approaches to factuality evaluation. Section 4 then shows how to integrate the function directly into the formal semantics of a paraconsistent logic through (ii) an LLM-grounded interpretation that preserves the soundness and completeness of analytic tableau systems for reasoning in the logic. Section 5 provides (iii) empirical evidence that the approach can be practically implemented, presenting findings from an evaluation using benchmarks derived from short-form factuality benchmarks and discussing limitations with potential mitigations. Section 6 concludes with a summary and discussion of future work.
@@ 2 Related work
Current approaches to reasoning with LLMs (Hoppe et al., 2025; Cheng et al., 2025) are illustrated in Figure 2. In a prompt-based approach (a), an LLM is prompted with a verbalization $\delta(\Gamma)\in\Sigma^{*}$ of a set of formulas $\Gamma$, and performs natural language reasoning to produce a verbalization $\delta(\varphi)$ of $\varphi$ (Wei et al., 2022; Kojima et al., 2022; Dhuliawala et al., 2023; Yao et al., 2023). In a solver-based approach (b), an LLM is prompted with a verbalization of a set of formulas and produces a set of formulas $\Gamma$ in an object language $\mathcal{L}$, which a reasoner uses to deduce $\varphi$ (Pan et al., 2023; Olausson et al., 2023; Callewaert et al., 2025). In an approach based on pre-training or fine-tuning (c), a reasoner provides a training set of proofs $\Pi$, and the LLM learns from that to reason as in (a) (Jiao et al., 2023; Morishita et al., 2024; Feng et al., 2024; Liu et al., 2025). Unlike approaches (b) and (c) that use LLMs alongside reasoning systems, in our proposed interpretation-based approach (d) we integrate LLMs directly into the formal semantics of the reasoner’s logic itself, by using an LLM to implement an interpretation function $\mathcal{I}$ to be used by the reasoner in inferring $\varphi$. This allows us to provide formal guarantees about the soundness and completeness of the reasoning process that incorporates the LLM.
@@
Figure 2: Four different approaches to logical reasoning with LLMs. Let $\mathcal{L}$ be a first-order language, $\Gamma$ be a set of statements in $\mathcal{L}$, $\varphi,\psi$ be statements in $\mathcal{L}$, $\mathcal{I}$ be an interpretation for $\mathcal{L}$, and $\delta:\mathcal{P}(\mathcal{L})\rightarrow\Sigma^{*}$ be a verbalization function that takes a set of formulas in $\mathcal{L}$ and returns a natural language translation of the formulas. In each approach, we show how reasoning is performed in the context of generating a formal or natural language proof showing that $\Gamma\vdash\varphi$.
@@ 3 Bilateral factuality evaluation of an atomic formula using an LLM
First, we generate a natural language verbalization of an atomic formula. We then prompt the LLM to generate two statements as to the verifiability and refutability of the assertion represented by the verbalization. The statements are then mapped into the set of truth values $\mathcal{V}_{3}$ used in weak Kleene logic (Kleene, 1952; Szmuc, 2019), i.e., $\mathfrak{t}$ (true), $\mathfrak{e}$ (undefined), and $\mathfrak{f}$ (false). Weak Kleene truth values allow us to formalize the semantics of LLM-judge output. For example, the SimpleQA grader used in the SimpleQA benchmark (Wei et al., 2024) grades an LLM’s answer to a question as either “CORRECT”, “INCORRECT”, or “NOT ATTEMPTED”; the PreciseWikiQA Question Answerability Prompt in the HalluLens benchmark (Bang et al., 2025) uses “UNVERIFIABLE” instead of “NOT ATTEMPTED”. We equate “CORRECT” with $\mathfrak{t}$ and “INCORRECT” with $\mathfrak{f}$; equating $\mathfrak{e}$ with “NOT ATTEMPTED” or “UNVERIFIABLE” is consistent with Kleene’s original statement that it indicates “an absence of information” that a given formula is either $\mathfrak{t}$ or $\mathfrak{f}$ (Kleene, 1952, p. 333).
Finally, we pair the weak Kleene truth value $u$ for verifiability with the weak Kleene truth value $v$ for refutability to yield a generalized truth value $\langle u,v\rangle\in\mathcal{V}_{3}\times\mathcal{V}_{3}$. This type of generalized truth value offers significant advantages over truth valuations provided by current approaches to factuality evaluation using LLM judges: they enable systematic distinctions between different types of evidence (Shramko and Wansing, 2011; Ferguson, 2021b), provide principled methods for handling inconsistent and incomplete information (Shramko and Wansing, 2011; Szmuc, 2019; Ferguson, 2021b; Correia, 2010), and enhance the explanatory transparency of logical valuations through their structured nature (Shramko and Wansing, 2011; Ferguson, 2021b; Fine, 2016). Figure 3 shows an example of this process in action; a longer example is provided in Appendix C.
@@
In practice, $\delta$ can be implemented in a number of ways: for example, through a template-based approach (Ell and Harth, 2014), by prompting an LLM to generate a verbalization of the set of formulas (Perevalov and Both, 2024), or simply providing the formulas directly in the syntax of the given object language.
@@
The “otherwise” cases in Definition 3.4 reflect the possibility of LLMs failing to output tokens indicating the state of verification or refutation, for example, by failing to follow prompt instructions, or timing out during a call to an inference API. Additionally, in our implementation, we use repeated sampling with majority vote (Brown et al., 2024) to determine each of the truth value components; other approaches to hallucination mitigation, such as chain-of-verification (Dhuliawala et al., 2023), are admissible as well.
The definition of $\zeta$ leaves open the possibility that multiple calls over time might return different truth values. Analytic tableau reasoning assumes that atomic formulas are assigned a truth value that is stable within the scope of the reasoning process. We ensure this by using a caching version of the bilateral evaluation function $\zeta_{c}$ where valuations for atomic formulas are persistently and immutably stored. This type of caching is consistent with the range of optimization techniques used in description logics tableau reasoners (Goré and Nguyen, 2007; Nguyen, 2009).
@@ 4 LLM-grounded interpretations
Lemma 4.2 effectively states that the logical validity of derivations depends on the structure of the logical operators rather than the specific content of atomic propositions, which is a fundamental principle in formal logic. The tableau method only depends on the truth-functional behavior of the logical connectives, which remains unchanged between standard and LLM-grounded interpretations.
@@ 5 Evaluation
To validate the feasibility of LLM-grounded interpretations, we evaluate the bilateral evaluation function $\zeta$ that underlies them, using question/answer pairs from two short-form factuality benchmarks. Question/answer pairs in short-form factuality benchmarks are typically factoids providing a question together with a short answer (Figure 3). We use these as an approximation to the verbalization $\delta(\varphi)$ of an atomic formula $\varphi$. We used the short-form factuality benchmarks GPQA (Rein et al., 2023) and SimpleQA (Wei et al., 2024) to create two balanced test datasets (each N=400) for our experiments. Test data preparation and experimental setup are discussed in Appendix D. We evaluated the performance of $\zeta$ over the two datasets using two metrics: macro F1 against question/answer pairs that the judge did not abstain from, i.e., where the judge provided a valuation of $\zeta(\varphi)=\langle\mathfrak{t},\mathfrak{f}\rangle$ (i.e., verified and not refuted) or $\zeta(\varphi)=\langle\mathfrak{f},\mathfrak{t}\rangle$ (i.e., not verified and refuted), and coverage, which is the percentage of the total set of question/answer pairs where the judge did not abstain. We also measured the time taken per evaluation, and the number of tokens used per evaluation.
@@ Results
Dataset Model Type $𝔱\langle\mathfrak{t},\mathfrak{t}\rangle$,$\langle\mathfrak{t},\mathfrak{f}\rangle$ $\langle\mathfrak{f},\mathfrak{t}\rangle$ $\langle\mathfrak{f},\mathfrak{f}\rangle$
@@ 6 Conclusion and future work
In conclusion, we have described a novel approach to logical reasoning using LLMs, with several key contributions. We defined a bilateral approach to factuality evaluation that allows us to identify gaps and contradictions in an LLM’s parametric knowledge. We introduced the concept of LLM-grounded interpretations that integrate an LLM directly into the formal semantics of the underlying logic while preserving its soundness and completeness. We provided empirical evidence that this formal framework can be realized in practice, providing a path to enabling LLMs to serve as broad-coverage knowledge sources for logical reasoners.
@@ A.2 Truth functions
˙$\dot{\exists}(X)$ =$=\begin{cases}\mathfrak{t}&\text{if }\langle\mathfrak{t},\mathfrak{t}\rangle\in X\\ \mathfrak{e}&\text{if for all }\langle u\text{, }v\rangle\text{, either }u=\mathfrak{e}\text{ or }v=\mathfrak{e}\\ \mathfrak{f}&\text{if }\langle\mathfrak{t},\mathfrak{t}\rangle\notin X\text{ and for some }\langle u,v\rangle\in X\text{, }u\neq\mathfrak{e}\text{ and }v\neq\mathfrak{e}\end{cases}$
˙$\dot{\forall}(X)$ =$=\begin{cases}\mathfrak{t}&\text{if }\langle\mathfrak{t},\mathfrak{f}\rangle\text{, }\langle\mathfrak{t},\mathfrak{e}\rangle\notin X\text{ and for some }\langle u\text{, }v\rangle\in X,u\neq\mathfrak{e}\text{ and }v\neq\mathfrak{e}\\ \mathfrak{e}&\text{if for all }\langle u\text{, }v\rangle\in X\text{, either }u=\mathfrak{e}\text{ or }v=\mathfrak{e}\\ \mathfrak{f}&\text{if }\{\langle\mathfrak{t},\mathfrak{t}\rangle\text{, }\langle\mathfrak{t}\text{, }\mathfrak{e}\rangle\}\cap X\neq\emptyset\text{ andfor some }\langle u,v\rangle\in X\text{, either }u=\mathfrak{e}\text{ or }v=\mathfrak{e}\end{cases}$
@@ D.4 Truth value distributions
Judge Model Prompt $𝔱\langle\mathfrak{t},\mathfrak{t}\rangle$,$\langle\mathfrak{t},\mathfrak{f}\rangle$ $\langle\mathfrak{f},\mathfrak{t}\rangle$ $\langle\mathfrak{f},\mathfrak{f}\rangle$
@@
Judge Model Prompt $𝔱\langle\mathfrak{t},\mathfrak{t}\rangle$,$\langle\mathfrak{t},\mathfrak{f}\rangle$ $\langle\mathfrak{f},\mathfrak{t}\rangle$ $\langle\mathfrak{f},\mathfrak{f}\rangle$

### v2

# Sound and Complete Neurosymbolic Reasoning with LLM-Grounded Interpretations

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in natural language understanding and generation, but they exhibit problems with logical consistency in the output they generate. How can we harness LLMs’ broad-coverage parametric knowledge in formal reasoning despite their inconsistency? We present a method for directly integrating an LLM into the interpretation function of the formal semantics for a paraconsistent logic. We provide experimental evidence for the feasibility of the method by evaluating the function using datasets created from several short-form factuality benchmarks. Unlike prior work, our method offers a theoretical framework for neurosymbolic reasoning that leverages an LLM’s knowledge while preserving the underlying logic’s soundness and completeness properties.
@@ 1 Introduction
Applications involving commonsense reasoning and biomedical knowledge, where inconsistencies and incomplete information are commonplace, remain challenging frontiers for AI systems. While LLMs encode vast parametric knowledge (Petroni et al., 2019), they also suffer from inconsistency and incompleteness (Cheng et al., 2025), limiting their use as knowledge bases for such applications. Efforts to connect logical reasoning with LLMs relying on prompting strategies and external symbolic solvers (Wei et al., 2022; Cheng et al., 2025) show potential but exhibit significant shortcomings as well Hoppe et al. (2025), lacking formal frameworks for managing LLM knowledge inconsistency and incompleteness. Paraconsistent logics (Priest et al., 2025) are multi-valued non-classical logics that handle inconsistent information without logical explosion, where contradictions would make everything provable. Belnap computers (Belnap, 1977a, b) are theoretical constructions described by Nuel Belnap to model contexts in which machines are responsible for reasoning in the face of incomplete or inconsistent information. This raises a natural question: can paraconsistent logic and Belnap computers provide a more natural framework for using LLMs as knowledge bases, despite their inconsistency and incompleteness?
We propose a Belnap computer using an LLM judge as an external knowledge source. LLM judges scale factuality evaluation of LLM output for short- or long-form question answering tasks, returning truth valuations for statements in the LLM’s output (Li et al., 2024). In our approach, an LLM judge responds to a query for the valuation of an atomic formula in the context of a paraconsistent reasoner with a generalized truth value (Shramko and Wansing, 2025, 2011). Generalized truth values allow the LLM judge to provide information to the reasoner not only about the degree of truth of an atomic formula given an LLM’s parametric knowledge, but also with respect to the degree of knowledge the LLM has about the formula.
@@
Section 2 discusses related work, and Section 3 defines (i) a bilateral factuality evaluation function that provides information beyond current LLM judge approaches to factuality evaluation. Section 4 integrates the function directly into the formal semantics of a paraconsistent logic through (ii) an LLM-grounded interpretation that preserves the soundness and completeness of analytic tableau systems for reasoning in the logic. Section 5 provides (iii) empirical evidence for practical implementation, presenting evaluation findings using benchmarks derived from short-form factuality benchmarks and discussing limitations. Section 6 concludes with a summary and discussion of future work.
@@ 2 Related work
Current approaches to reasoning with LLMs (Hoppe et al., 2025; Cheng et al., 2025) appear in Figure 2. In a prompt-based approach (a), an LLM is prompted with a verbalization $\delta(\Gamma)\in\Sigma^{*}$ of a set of formulas $\Gamma$, and performs natural language reasoning to produce a verbalization $\delta(\varphi)$ of $\varphi$ (Wei et al., 2022; Kojima et al., 2022; Dhuliawala et al., 2023; Yao et al., 2023). In a solver-based approach (b), an LLM is prompted with a verbalization of a set of formulas and produces a set of formulas $\Gamma$ in an object language $\mathcal{L}$, which a reasoner uses to deduce $\varphi$ (Pan et al., 2023; Olausson et al., 2023; Callewaert et al., 2025). In an approach based on pre-training or fine-tuning (c), a reasoner provides a training set of proofs $\Pi$, and the LLM learns from that to reason as in (a) (Jiao et al., 2023; Morishita et al., 2024; Feng et al., 2024; Liu et al., 2025). Unlike approaches (b) and (c) that use LLMs alongside reasoning systems, in our proposed interpretation-based approach (d) we integrate LLMs directly into the formal semantics of the reasoner’s logic itself, by using an LLM to implement an interpretation function $\mathcal{I}$ to be used by the reasoner in inferring $\varphi$. This allows us to provide formal guarantees about the soundness and completeness of the reasoning process that incorporates the LLM. In contrast with approaches (a), (b), and (c), instead of trying to get an LLM to reason using logic, we get a logic to reason using an LLM.
@@
Figure 2: Four different approaches to logical reasoning with LLMs. Let $\mathcal{L}$ be a first-order language, $\Gamma$ be a set of statements in $\mathcal{L}$, $\varphi$ be a statement in $\mathcal{L}$, $\mathcal{I}$ be an interpretation for $\mathcal{L}$, $\Pi$ be a set of proofs of statements in $\mathcal{L}$, and $\delta:\mathcal{P}(\mathcal{L})\rightarrow\Sigma^{*}$ be a verbalization function that takes a set of formulas in $\mathcal{L}$ and returns a natural language translation of the formulas. In each approach, we show how reasoning is performed in the context of generating a formal or natural language proof showing that $\Gamma\vdash\varphi$.
@@ 3 Bilateral factuality evaluation of an atomic formula using an LLM
First, we generate a natural language verbalization of an atomic formula, then prompt the LLM to generate two statements on the verifiability and refutability of the assertion. The statements are then mapped into the set of truth values $\mathcal{V}_{3}$ used in weak Kleene logic (Kleene, 1952; Szmuc, 2019), i.e., $\mathfrak{t}$ (true), $\mathfrak{e}$ (undefined), and $\mathfrak{f}$ (false). Weak Kleene truth values allow us to formalize the semantics of LLM-judge output. For example, the SimpleQA grader used in the SimpleQA benchmark (Wei et al., 2024) grades an LLM’s answer to a question as either “CORRECT”, “INCORRECT”, or “NOT ATTEMPTED”; the PreciseWikiQA Question Answerability Prompt in the HalluLens benchmark (Bang et al., 2025) uses “UNVERIFIABLE” instead of “NOT ATTEMPTED”. We equate “CORRECT” with $\mathfrak{t}$ and “INCORRECT” with $\mathfrak{f}$; equating $\mathfrak{e}$ with “NOT ATTEMPTED” or “UNVERIFIABLE” is consistent with Kleene’s original statement that it indicates “an absence of information” that a given formula is either $\mathfrak{t}$ or $\mathfrak{f}$ (Kleene, 1952, p. 333).
Finally, we pair the weak Kleene truth value $u$ for verifiability with the weak Kleene truth value $v$ for refutability to yield a generalized truth value $\langle u,v\rangle\in\mathcal{V}_{3}\times\mathcal{V}_{3}$. Generalized truth values offer significant advantages over truth values provided by current approaches to factuality evaluation using LLM judges: they enable systematic distinctions between different types of evidence (Shramko and Wansing, 2011; Ferguson, 2021b), provide principled methods for handling inconsistent and incomplete information (Shramko and Wansing, 2011; Szmuc, 2019; Ferguson, 2021b; Correia, 2010), and enhance the explanatory transparency of logical valuations through their structured nature (Shramko and Wansing, 2011; Ferguson, 2021b; Fine, 2016). Figure 3 shows an example of this process in action; a longer example is provided in Appendix C.
@@
In practice, $\delta$ can be implemented through a template-based approach (Ell and Harth, 2014), by prompting an LLM to generate a verbalization (Perevalov and Both, 2024), or by providing formulas directly in the object language syntax.
@@
The “otherwise” cases in Definition 3.4 reflect LLMs failing to output tokens indicating verification or refutation state, e.g., by failing to follow instructions or timing out during API calls. We use repeated sampling with majority vote (Brown et al., 2024) to determine truth value components; other hallucination mitigation approaches, such as chain-of-verification (Dhuliawala et al., 2023), are also admissible.
The definition of $\zeta$ leaves open the possibility that multiple calls over time might return different truth values. The definition of $\zeta$ allows multiple calls to return different truth values. Analytic tableau reasoning assumes atomic formulas have stable truth values within the reasoning process scope. We ensure this by using a caching version of the bilateral evaluation function $\zeta_{c}$ where valuations for atomic formulas are persistently and immutably stored. This type of caching is consistent with the range of optimization techniques used in description logics tableau reasoners (Goré and Nguyen, 2007; Nguyen, 2009).
@@ 4 LLM-grounded interpretations
Lemma 4.2 states that logical validity of derivations depends on logical operator structure rather than atomic proposition content, a fundamental principle in formal logic. The tableau method only depends on the truth-functional behavior of the logical connectives, which remains unchanged between standard and LLM-grounded interpretations.
@@ 5 Evaluation
[…]
@@ Results
Dataset Model Type $\langle\mathfrak{t},\mathfrak{t}\rangle$ $\langle\mathfrak{t},\mathfrak{f}\rangle$ $\langle\mathfrak{f},\mathfrak{t}\rangle$ $\langle\mathfrak{f},\mathfrak{f}\rangle$
@@ 6 Conclusion and future work
We described a novel approach to logical reasoning using LLMs with several key contributions. We defined a bilateral approach to factuality evaluation that identifies gaps and contradictions in LLM parametric knowledge. We introduced the concept of LLM-grounded interpretations that integrate an LLM directly into the formal semantics of the underlying logic while preserving its soundness and completeness. We provided empirical evidence that this formal framework can be realized in practice, providing a path to enabling LLMs to serve as broad-coverage knowledge sources for logical reasoners.
@@ Acknowledgements
This work was partially supported by the EU’s Horizon Europe research and innovation programme within the ENEXA project (grant Agreement no. 101070305). Particular thanks are due to Fabian Hoppe, Levin Hornischer, Jan-Christoph Kalo, and Lise Stork for discussions and perceptive observations that enriched our research.
@@ A.2 Truth functions
[…]
@@ D.4 Truth value distributions
Judge Model Prompt $\langle\mathfrak{t},\mathfrak{t}\rangle$ $\langle\mathfrak{t},\mathfrak{f}\rangle$ $\langle\mathfrak{f},\mathfrak{t}\rangle$ $\langle\mathfrak{f},\mathfrak{f}\rangle$
@@
Judge Model Prompt $\langle\mathfrak{t},\mathfrak{t}\rangle$ $\langle\mathfrak{t},\mathfrak{f}\rangle$ $\langle\mathfrak{f},\mathfrak{t}\rangle$ $\langle\mathfrak{f},\mathfrak{f}\rangle$